OpenStudy (anonymous):

explain how to simplify x(2x+3x^2) Please

OpenStudy (zzr0ck3r):

distribute the x to both terms in the parenthesis.

OpenStudy (anonymous):

2x+3x^3

OpenStudy (anonymous):

Multiply each of the terms by x, x(2x+3x^2) = x*2x + x*3x^2= 2x^2+3x^3

OpenStudy (anonymous):

yeah opps i made a mistake. @Dahlioz is correct

OpenStudy (anonymous):

thanks people

OpenStudy (anonymous):

np